OpenStudy (anonymous):

Hi guys. I have been using this website and it is such a tremendous help!!! The people on here are so resourceful and great at teaching so I figured I could put this question on here too and would get some help. So here it is: could anyone help with Properties of Logarithms Product Property Quotient Property Power Property I need and explanation and simple example problem for each. Please if you could help me out. Anything helps!

OpenStudy (bahrom7893):

Which one don't you understand?

OpenStudy (anonymous):

All of Them, I really need a quick lesson to learn them.

OpenStudy (anonymous):

Like I said, anything helps.

OpenStudy (bahrom7893):

Okay, let's start with the first one - what is the Product Property?

OpenStudy (bahrom7893):

Define it first. I can tell you the definition, but why don't you at least look it up and post it yourself?

OpenStudy (anonymous):

is it: is the sum or difference of logarithms and the number to a power is the power times the logarithm of that number. I think

OpenStudy (bahrom7893):

Okay, that can also be written as: log(A*B) = log(A) + log(B) log(A/B) = log(A) - log(B) log(A^b) = b*log(A)

OpenStudy (bahrom7893):

now just plug in some numbers instead of A and B and those will be your examples.

OpenStudy (anonymous):

Oh, um. Okay thank you.

OpenStudy (anonymous):

Are there equations for the other ones as well?

OpenStudy (bahrom7893):

those are all 3

OpenStudy (anonymous):

Oh! duuhhh. thanks